EXERCISE 11–1 Compute the Return on Investment (ROI) [LO11–1] Guided Example 11-1 Alyeska Services Company, a division of a major oil company, provides various services to the operators of the North Slope oil field in Alaska. Data concerning the most recent year appear below

Compute the margin for Alyeska Services Company.

Compute the turnover for Alyeska Services Company.

Compute the return on investment (ROI) for Alyeska Services Company.

Sales .......... Net operating income ............ Average operating assets ......... $7,500,000 $600,000 $5,000,000
Answer #1

Margin=Net operating income/Sales

=(600,000/7,500,000)=8%

Turnover=Sales/Average operating assets

=(7,500,000/5,000,000)=1.5

ROI=Net operating income/Average operating assets

=(600,000/5,000,000)=12%